OpenStudy (anonymous):

Simplify 9 to the 2nd over 9 to the 7th.

Nnesha (nnesha):

\(\Huge \color{goldenrod}{\star^{ \star^{\star:)}}}\Huge \color{green}{\star^{ \star^{\star:)}}}\)\(\Huge\color{salmon}{\scr Welcome ~to~OpenStudy!}\) \(\Huge \color{blue}{\star^{ \star^{\star:)}}}\Huge \color{red}{\star^{ \star^{\star:)}}}\) \(\Huge \color{navy}{\star^{ \star^{\star:)}}}\Huge \color{purple}{\star^{ \star^{\star:)}}}\) \(\Huge \color{aquamarine}{\star^{ \star^{\star:)}}}\Huge \color{orangered}{\star^{ \star^{\star:)}}}\)\(\Huge \color{teal}{\star^{ \star^{\star:)}}}\Huge \color{plum}{\star^{ \star^{\star:)}}}\) \(\Huge \color{seagreen}{\star^{ \star^{\star:)}}}\)

Nnesha (nnesha):

\[\huge\rm \frac{ 9^2 }{ 9^7 }\] when we divide same bases we should ` subtract` exponents \[\huge\rm \frac{ x^m}{ x^n }=x^{m-n}\]

OpenStudy (anonymous):

9^5

Nnesha (nnesha):

nope

OpenStudy (anonymous):

im not good at math

Nnesha (nnesha):

aww you will be an expert \[\huge\rm \frac{ 9^2 }{ 9^7 }=9^{2\color{red}{-}7}\] it should be 2-7

OpenStudy (anonymous):

you lost me im sorry

Nnesha (nnesha):

we should subtract top exponent from the bottom exponent so it should be 2-7 and \[2-7\cancel{=}5\] sign error!

OpenStudy (anonymous):

so it would be 9 to the 5th power right?

Nnesha (nnesha):

no 2-7 isn't equal to 5 remember when we subtract if bigger number is negative then answer would be negative !

Nnesha (nnesha):

2-3= -1

Nnesha (nnesha):

example^

OpenStudy (anonymous):

ohh yea im dumb

OpenStudy (anonymous):

it would be 9 to the -1

Nnesha (nnesha):

nope 2-7 = ???

OpenStudy (anonymous):

i mean -5

Nnesha (nnesha):

yes right now we need to convert negative to positive exponent so apply this exponent rule \[\huge\rm x^{-m}=\frac{ 1 }{ x^m } \]

Nnesha (nnesha):

so \[9^{-5}= ?\]

OpenStudy (anonymous):

1/9 to the -5th power omg thx you help alot can you syick around incace i need help again ;-)

Nnesha (nnesha):

when we flip the fraction sign would change

Nnesha (nnesha):

so 1/9 to the what power ?

OpenStudy (anonymous):

-5

Nnesha (nnesha):

nope look at this example \[\huge\rm x^{-m}=\frac{ 1 }{ x^m }\] we don't want negative exponent that's the reason we should flip the fraction

Nnesha (nnesha):

so change the sign when you flip it

OpenStudy (anonymous):

im really confused

Nnesha (nnesha):

i'll give you an example \[\large\rm 3^{-4}=\frac{ 1 }{ 3^4 }\]

OpenStudy (anonymous):

ohhhhh 5

Nnesha (nnesha):

yes right!

OpenStudy (anonymous):

thz you

Nnesha (nnesha):

np :=)

OpenStudy (anonymous):

Which expression is equivalent to (5^3)^−2? can u help me with this one plzzzzz

Nnesha (nnesha):

\[\huge\rm (x^m)^n=x^{m \times n}\] you just need to know exponents rules!

OpenStudy (anonymous):

5^-6

Nnesha (nnesha):

yes so what about negative e xponents ?

Nnesha (nnesha):

what would be ur next step ?

OpenStudy (anonymous):

5^6

OpenStudy (anonymous):

is that the anwser

Nnesha (nnesha):

nope

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha nope look at this example \[\huge\rm x^{-m}=\frac{ 1 }{ x^m }\] we don't want negative exponent that's the reason we should flip the fraction \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

this is frustrating idk how to do this

Nnesha (nnesha):

\[5^{-6}\] is same as \[9^{-5}\] so how did you changed 9^-5 to positive exponent

OpenStudy (anonymous):

would it be1/5^3 juss a ruf guess

Nnesha (nnesha):

remember 5 to the -6 power is same as \[\huge\rm \frac{ 5^{-6} }{ 1}\] now flip the fraction and change the sign of the exponent

Nnesha (nnesha):

it's 5 to the -6 power not 3

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @copper224 would it be1/5^3 juss a ruf guess \(\color{blue}{\text{End of Quote}}\) 5 to the what power ?

OpenStudy (anonymous):

6th im sorry if im frustrating you

OpenStudy (anonymous):

i juss suck at math

Nnesha (nnesha):

im fine :D

Nnesha (nnesha):

well you're in a learning process you will be good at it!

Nnesha (nnesha):

alright good luck! practice!!!

OpenStudy (anonymous):

In which expression should the exponents be multiplied? one fifth to the 2nd times one fifth to the 6th 9 to the 3rd over 9 to the 4th 73 ⋅ 78 (26)−5

OpenStudy (anonymous):

@nNesha